OpenStudy (anonymous):

order: Find the gradient of the normal to the curve, at the specified value of t. \[x=\cos^3t~~~~y=\sin^3t~~~~~~when ~~~t={1\over6}\pi\]

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

x'= -3cos^2tsint??? y'= 3sin^2tcost???

OpenStudy (dumbcow):

yes :)

OpenStudy (anonymous):

\[3\sin^2tcost \over -3\cos^2tsint\]

OpenStudy (anonymous):

\[-sint \over cost\]?

OpenStudy (dumbcow):

good

OpenStudy (anonymous):

-tant?

OpenStudy (dumbcow):

yes either way would work since you still have to evaluate at t=pi/6

OpenStudy (anonymous):

Ok. let me evaluate. How can you do this using pi ? I always convert to decimals...?

OpenStudy (dumbcow):

pi/6 converted to degrees is 180/6 or 30

OpenStudy (anonymous):

OK -tan30= -0.57735 m(-0.57735)=-1 m=1.73205 How do I convert this into sqrt?

OpenStudy (dumbcow):

looks good to leave it in radicals -tan(30) = -1/sqrt3 m = sqrt3

OpenStudy (anonymous):

Is there some sort of formula for radicals?

OpenStudy (dumbcow):

no not really, well the unit circle since 30 degrees is on there sin(30) = 1/2 cos(30) = sqrt3/2 -tan(30) = -1/2 * 2/sqrt3 = -1/sqrt3

OpenStudy (anonymous):

Hm Ok. Thank you! In exams, do they regard radicals over decimals? Or do you not know?

OpenStudy (dumbcow):

depends on the teacher or policies of school radical is considered more exact while decimal are approximate answers

OpenStudy (anonymous):

Hm Ok. It's external examinations.

OpenStudy (anonymous):

is this trigonometry? @dum

OpenStudy (anonymous):

@dumbcow

OpenStudy (anonymous):

Trig, yes and parametric equations

OpenStudy (dumbcow):

yeah ^^

OpenStudy (anonymous):

ooh, more advance,.

OpenStudy (anonymous):

Thanks for helping!